OpenStudy (anonymous):

I am really having a hard time in graphing this equation y = -3x+4

OpenStudy (anonymous):

solve for the y and x intercepts

OpenStudy (anonymous):

When X is 0, then Y is 4 (y-intercept). From there, just plug numbers in for x and graph each point. -2 = 16 -1 = 7 0 = 4 1 = 1 2 = -2 Do you see what I am doing?

OpenStudy (anonymous):

You can do it with any numbers you like. You can also solve for y for the x intercept. Make y = 0 and solve. \[0=-3x+4\] Subtract 4 from each side \[-4=-3x\] Divide both sides by -3 \[\frac{ -4 }{ -3 }=\frac{ 4 }{ 3 }=x\] So 4/3 is your x intercept.

OpenStudy (anonymous):

now that youalrady have your x and y intercepts w/c are: x-int: (4/3, 0) y-int: (0, 4) follow frizzy's first comment: